find the vertex of y=x^2-1
Answer by stanbon(75887)

find the vertex of y=x^2-1
a = 1 ; b = 0 ; c = -1
----
vertex occurs at x = -b/(2a) = -0/(2) = 0
f(0) = -1
----
Vertex: (0,-1)
=======================
Cheers,
Stan H.
==================